OpenStudy (anonymous):

Tell whether the function f(x) = -x3 + 7x + 4 is an even or odd function, and describe its symmetry.

OpenStudy (anonymous):

for this you need to check f(x) and f(-x) f(x)=-x^3+7x+4 f(-x)=x^3-7x+4 they are even if f(x)=f(-x) odd if f(x)=-f(x) here none of this is true, so the function is not even or odd

OpenStudy (anonymous):

your right